If the roots of the cubic az3 + bz2 + cz + d = 0 form an arithmetic progression α − β,
α, α + β, prove that (2b2 − 9ac)b + 27a2d = 0.
1
Expert's answer
2020-05-19T08:48:03-0400

We are given that "az^3 + bz^2 + cz + d = 0" has three roots in arithmetic progression "\u03b1 \u2212 \u03b2,\n\n\u03b1, \u03b1 + \u03b2".

Now, we known as: Sum of roots = "-b\/a" , Product of roots = "-d\/a" and sum of product of two roots = "c\/a".

So, we got three equations in our problem as:

"(\alpha−\beta)+(\alpha)+(\alpha+\beta)= -\frac{b}{a} \\ 
(\alpha−\beta)(\alpha)(\alpha+\beta) = -\frac{d}{a} \\
(\alpha−\beta)(\alpha)+ (\alpha)(\alpha+\beta)+ (\alpha−\beta)(\alpha+\beta) = \frac{c}{a}"

So, "3\alpha = -\frac{b}{a}, \alpha(\alpha^2-\beta^2) = -\frac{d}{a}" and "3\alpha^2-\beta^2 = \frac{c}{a}" .


So by first and last equation we got, "\alpha = -\frac{b}{3a}" and "\beta^2 = \frac{b^2}{3a^2}-\frac{c}{a}".


Now from middle equation we get,

"(-\frac{b}{3a})(\frac{b^2}{9a^2} - \frac{b^2}{3a^2} + \frac{c}{a} ) = -\frac{d}{a}"

"\implies \frac{b(b^2 - 3b^2 + 9ac )}{9a^2} = 3d \\
\implies (2b^2 - 9ac)b + 27a^2d = 0"